OpenStudy (anonymous):

solvw 3sqr rt symbol x+5 inside the symbol = 4

OpenStudy (anonymous):

\[3\sqrt{x+5}=4\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

9(x+5)=16 x+5=16/9 x=16/9-5 x=(16-45)/9 x=-29/9

OpenStudy (anonymous):

incorrect

OpenStudy (anonymous):

\[\sqrt[3]{x+5}=4\]

OpenStudy (anonymous):

o ok then x+5=4^3 x+5=64 x=59